# Generalizing Denoising to Non-Equilibrium Structures Improves Equivariant Force Fields

### v1

@@ Results.
We train Equiformer ($L_{max}=2$(Liao & Smidt, 2023) with DeNS based on their implementation and summarize the results in Table 4. DeNS improves the results on all molecules, and Equiformer ($L_{max}=3$) trained with DeNS achieves overall best results. Particularly, Equiformer ($L_{max}=2$) trained with DeNS acheives better results on all the tasks and requires 3.1×3.1$\times$3.1 × less training time than Equiformer ($L_{max}=3$) trained without DeNS. This demonstrates that for this small dataset, training an auxiliary task and using data augmentation are more efficient and result in larger performance gain than increasing $L_{max}$ from $2$ to $3$. Additionally, we find that the gains from training DeNS as an auxiliary task are comparable to pre-training. Zaidi et al. (2023) uses TorchMD-NET (Thölke & Fabritiis, 2022) pre-trained on the PCQM4Mv2 dataset and report results on Aspirin. Their improvement on force MAE is about $17.2\%$ (Table 3 in Zaidi et al. (2023)). Training Equiformer with DeNS results in $20.8\%$ improvement on force MAE without relying on another dataset. Note that we only increase training time by $10.5\%$ while their method takes much more time since PCQM4Mv2 dataset is more than 3000×3000$\times$3000 × larger than MD17. Moreover, we also train variants of Equiformer ($L_{max}=2$) by removing attention and layer normalization to investigate the performance gain of DeNS on different network archtiectures. The results are summarized in Table 5, and DeNS improves all the models. We note that Equiformer without attention and layer normalization reduces to SEGNN (Brandstetter et al., 2022) but with a better radial basis function. Since the models cover many variants of equivariant networks, this suggests that DeNS is general and can be helpful to many equivariant networks.
Comparison between DeNS with and without Force Encoding.
@@
We use Equiformer ($L_{max}=2$) to compare DeNS with and without force encoding and demonstrate that force encoding is critical. The results are summarized in Table 6. DeNS with force encoding (Index 3) achieves the best results across all molecules. Compared to training without denoising (Index 1), DeNS without force encoding (Index 2) only achieves slightly better results on some molecules (i.e., Benzene, Malondaldehyde, and Toluene) and much worse results on others. For molecules on which DeNS without force encoding is helpful, adding force encoding can achieve even better results. For others, force encoding is indispensable for DeNS to be effective.
@@
Aspirin Benzene Ethanol Malonaldehyde Naphthalene Salicylic acid Toluene Uracil
Index Method energy forces energy forces energy forces energy forces energy forces energy forces energy forces energy forces
1 Equiformer ($L_{max}=2$) 5.3 7.2 2.2 6.6 2.2 3.1 3.3 5.8 3.7 2.1 4.5 4.1 3.8 2.1 4.3 3.3
2 Equiformer ($L_{max}=2$) + DeNS without force encoding 8.6 9.1 2.3 6.3 2.3 3.3 3.2 5.8 7.7 6.1 5.2 10.6 3.7 2.0 5.5 6.5
3 Equiformer ($L_{max}=2$) + DeNS with force encoding 5.1 5.7 2.3 6.1 2.2 2.6 3.2 4.4 3.7 1.7 4.3 3.9 3.5 1.9 4.2 3.3
@@
Table 6: Comparison between DeNS with and without force encoding on MD17 dataset. Mean absolute error results are evaluated on the MD17 testing set. Energy and force are in units of meV and meV/Å. Index 1 and Index 3 correspond to “Equiformer ($L_{max}=2$)” and “Equiformer ($L_{max}=2$) + DeNS” in Table 4.
@@ 5 Conclusion
In this paper, we propose to use denoising non-equilibrium structures (DeNS) as an auxiliary task to better leverage training data and improve performance of original tasks of energy and force predictions. Denoising non-equilibrium structures can be an ill-posed problem since there are many possible target structures. To address the issue, we propose to take the forces of original structures as inputs to specify which non-equilibrium structures we are denoising. With force encoding, DeNS successfully improves the performance of original tasks when it is used as an auxiliary task. We conduct extensive experiments on OC20, OC22 and MD17 datasets to demonstrate that DeNS can boost the performance of energy and force predictions with minimal increase in training cost and is applicable to many equivariant networks.
Acknowledgement
We thank Larry Zitnick, Xinlei Chen, Zachary Ulissi, Muhammed Shuaibi, Saro Passaro, Anuroop Sriram, and Brandon Wood for helpful discussions. We acknowledge the MIT SuperCloud and Lincoln Laboratory Supercomputing Center (Reuther et al., 2018) for providing high performance computing and consultation resources that have contributed to the research results reported within this paper.
Yi-Lun Liao and Tess Smidt were supported by DOE ICDI grant DE-SC0022215.
@@ Appendix A Related Works
A.1 SE(3)/E(3)-Equivariant Networks
@@
A.2 Comparison to previous works on denoising
We discuss the comparisons between previous works on denoising (Godwin et al., 2022; Zaidi et al., 2023; Feng et al., 2023b; Wang et al., 2023) and this work in chronological order as below.
Godwin et al. (2022) first proposes the idea of adding noise to 3D coordinates and then using denoising as an auxiliary task. The auxiliary task is trained along with the original task without relying on another large dataset. Their approach requires known equilibrium structures and therefore is limited to QM9 (Ramakrishnan et al., 2014; Ruddigkeit et al., 2012) and OC20 IS2RE datasets and can not be applied to force prediction such as OC20 S2EF dataset. For QM9, all the structures are at equilibrium, and for OC20 IS2RE, the target of denoising is the relaxed, equilibrium structure. Denoising without force encoding is well-defined on both QM9 and OC20 IS2RE. In contrast, this work proposes using force encoding to generalize their approach to non-equilibrium structures, which have much larger datasets than equilibrium ones. Force encoding can achieve better results on OC20 S2EF-2M dataset without any overhead (Index 1 and Index 2 in Table 1(e)) and is indispensable on MD17 dataset (Section 4.3).
Feng et al. (2023b) follows the same practice of pre-training via denoising (Zaidi et al., 2023) and proposes a different manner of adding noise. Specifically, they separate noise into dihedral angle noise and coordinate noise and only learn to predict coordinate noise. However, adding noise to dihedral angles requires tools like RDKit to obtain rotatable bonds and cannot be applied to other datasets like OC20 and OC22.
Although Zaidi et al. (2023) and Feng et al. (2023b) report results of force prediction on MD17 dataset, they first pre-train models on PCQM4Mv2 dataset (Nakata & Shimazaki, 2017) and then fine-tune the pre-trained models on MD17 dataset. We note that their setting is different from ours since we do not use any dataset for pre-training. As for fine-tuning on MD17 dataset, Zaidi et al. (2023) simply follows the same practice in standard supervised training. Feng et al. (2023b) explores fine-tuning with objectives similar to Noisy Nodes (Godwin et al., 2022), but the performance gain is much smaller than ours. Concretely, in Table 5 in Feng et al. (2023b), the improvement on force prediction on Aspirin is about $2.6\%$ while we improve force MAE by $20.8\%$.
@@ B.1 Training Details
Since each structure in OC20 S2EF dataset has a pre-defined set of fixed and free atoms, when training DeNS, we only add noise to free atoms and denoise corrupted free atoms. Noise is added to all the free atoms, and therefore the corruption ratio $r_{\text{DeNS\leavevmode\nobreak\ }}$ is $1$. Additionally, we find that multi-scale noise works better than using a fixed $\sigma$ and thus corrupt structures with multi-scale noise when training DeNS.
We add an additional block of equivariant graph attention to EquiformerV2 for noise prediction. We mainly follow the hyper-parameters of training EquiformerV2 without DeNS on OC20 S2EF-2M and S2EF-All+MD datasets. For training EquiformerV2 on OC20 S2EF-All+MD dataset, we increase the number of epochs from $1$ to $2$ for better performance. This results in higher training time than other methods. However, we note that we already demonstrate training with DeNS can achieve better results given the same amount of training time in Table LABEL:tab:comparison_with_without_dens. Table 7 summarizes the hyper-parameters of training EquiformerV2 with DeNS for the ablation studies on OC20 S2EF-2M dataset in Section 4.1.1 and for the main results on OC20 S2EF-All+MD dataset in Section 4.1.2. Please refer to the work of EquiformerV2 (Liao et al., 2023) for details of the architecture.
V100 GPUs with 32GB are used to train models. We use $16$ GPUs for training on OC20 S2EF-2M dataset and $128$ GPUs for OC20 S2EF-All+MD dataset. The training time and the numbers of parameters of different models on OC20 S2EF-2M dataset can be found in Table LABEL:tab:comparison_with_without_dens. For OC20 S2EF-All+MD dataset, the training time is $65465$ GPU-hours and the number of parameters is $160$M.
@@ C.1 Training Details
Different from OC20, all the atoms in a structure in OC22 are free, and therefore we cannot limit denoising to a pre-defined set of atoms. Instead, we use partially corrupted structures as discussed in Section 3.2.3 and add noise to and denoise random subsets of atoms. We follow the practice on OC20 dataset and use multi-scale noise to corrupt structures.
We add an additional block of equivariant graph attention to EquiformerV2 for noise prediction. Table 8 summarizes the hyper-parameters for the results on OC22 in Table 3.
We use $32$ V100 GPUs (32GB) for training. The training time is $5082$ GPU-hours, and the number of parameters $127$M.
@@
Hyper-parameters S2EF-2M S2EF-All+MD
@@
Batch size $64$ $512$
@@
Upper bound on standard deviations of Gaussian noise $\sigma_{\text{high}}$ $0.75$
Probability of optimizing DeNS $p_{\text{DeNS\leavevmode\nobreak\ }}$ $0.25$
Loss coefficient for DeNS $\lambda_{\text{DeNS\leavevmode\nobreak\ }}$ $10$
Corruption ratio $r_{\text{DeNS}}$ $1.0$ $1.0$
Table 7: Hyper-parameters of training EquiformerV2 with DeNS on OC20 S2EF-2M dataset and OC20 S2EF-All+MD dataset. For OC20 S2EF-2M dataset, the DeNS-related hyper-parameters, $\sigma_{\text{high}}$, $p_{\text{DeNS}}$ and $\lambda_{\text{DeNS}}$, can be found in Table 1.
@@
Upper bound on standard deviations of Gaussian noise $\sigma_{\text{high}}$ $0.5$
@@
Loss coefficient for DeNS $\lambda_{\text{DeNS\leavevmode\nobreak\ }}$ $100$
@@ D.2 Training Details
We use the same codebase as Equiformer (Liao & Smidt, 2023) for experiments on the MD17 dataset and follow most of the original hyper-parameters for training with DeNS. For training DeNS, we use an additional block of equivariant graph attention for noise prediction, which slightly increases training time and the number of parameters. We use single-scale noise with a fixed standard deviation $\sigma$ when corrupting structures. The hyper-parameters introduced by training DeNS and the values of energy coefficient $\lambda_{E}$ and force coefficient $\lambda_{F}$ on different molecules can be found in Table 9. Empirically, we find that linearly decaying DeNS coefficient $\lambda_{\text{DeNS}}$ to 00 thoughout the training can result in better performance. For the Equiformer variant without attention and layer normalization, we find that using normal distributions to initialize weights can result in training divergence and therefore we use uniform distributions. For some molecules, we find training Equiformer variant without attention and layer normalization with DeNS is unstable and therefore reduce the learning rate to $3\times 10^{-4}$.
@@
Hyper-parameter Aspirin Benzene Ethanol Malonaldehyde Naphthalene Salicylic acid Toluene Uracil
Energy coefficient $\lambda_{E}$ 1 1 1 1 2 1 1 1
Force coefficient $\lambda_{F}$ 80 80 80 100 20 80 80 20
Probability of optimizing DeNS $p_{\text{DeNS}}$ 0.25 0.25 0.25 0.25 0.25 0.125 0.125 0.25
Standard deviation of Gaussian noises $\sigma$ 0.05 0.05 0.05 0.05 0.05 0.025 0.025 0.05
DeNS coefficient $\lambda_{\text{DeNS}}$ 5 5 5 5 5 5 5 5
Corruption ratio $r_{\text{DeNS}}$ 0.25 0.25 0.25 0.25 0.25 0.25 0.25 0.25
Table 9: Hyper-parameters of training with DeNS on the MD17 dataset. The other hyper-parameters not listed here are the same as the original Equiformer ($L_{max}=2$) trained without DeNS.

### v2

@@ 4.3 MD17 Dataset
We train Equiformer ($L_{max}=2$(Liao & Smidt, 2023) and Equiformer ($L_{max}=3$) with DeNS based on their official implementation, where $L_{max}$ denotes the maximum degree of equivariant representations. As shown in Table 4, DeNS improves the results on all molecules. Along with the results on OC20 and OC22 datasets, DeNS can generally improve the performance on force predictions with both direct and gradient methods. Particularly, Equiformer ($L_{max}=2$) trained with DeNS acheives better results on all the tasks and requires 3.1×3.1$\times$3.1 × less training time than Equiformer ($L_{max}=3$) trained without DeNS. This demonstrates that for this small dataset, training an auxiliary task and using data augmentation are more efficient and result in larger performance gain than increasing $L_{max}$ from $2$ to $3$. Additionally, we find that the gains from training DeNS as an auxiliary task are comparable to pre-training. For example, Zaidi et al. (2023) uses TorchMD-NET (Thölke & Fabritiis, 2022) pre-trained on the PCQM4Mv2 dataset and reports results on Aspirin. Their improvement in force MAE is about $17.2\%$ (Table 3 in Zaidi et al. (2023)). Training Equiformer ($L_{max}=2$) with DeNS results in $20.8\%$ improvement in force MAE without relying on another dataset. Note that we only increase training time by $10.5\%$ while their method takes much more time since PCQM4Mv2 dataset is more than 3000×3000$\times$3000 × larger than the training set of MD17. Moreover, training with DeNS enables consistent improvement in all molecules when increasing $L_{max}$ from $2$ to $3$, and Equiformer ($L_{max}=3$) trained with DeNS achieves overall best results. In contrast, when DeNS is not used, increasing $L_{max}$ from $2$ to $3$ can lead to overfitting and worse results on some molecules (i.e., Benzene and Uracil).
Effect of DeNS on Different Network Architectures.
We train variants of Equiformer ($L_{max}=2$) by removing attention and layer normalization to investigate the performance gain of DeNS on different network archtiectures. The results are summarized in Table 5, and DeNS improves all the model variants. We note that Equiformer without attention and layer normalization reduces to SEGNN (Brandstetter et al., 2022) but with a better radial basis function. Since the models cover many variants of equivariant networks, this suggests that DeNS is general and can be helpful to many equivariant networks.
Ablation Study on the Design Choices of DeNS.
@@
We use Equiformer ($L_{max}=2$) to justify the design choices of DeNS. The results are summarized in Table 6 and are similar to those on OC20 S2EF-2M dataset. Comparing Index 2 and Index 3, force encoding consistently results in significant improvement in energy and force MAE. Compared to training without denoising (Index 1), DeNS without force encoding (Index 3) only achieves slightly better results on some molecules (i.e., Benzene, Malondaldehyde, and Toluene) and much worse results on others. For molecules on which DeNS without force encoding is helpful, adding force encoding can achieve even better results. For others, force encoding is indispensable for DeNS to be effective. Comparing Index 2 and Index 4, predicting energy given corrupted structures results in overall better performance. Finally, the comparison between Index 2 and Index 5 demonstrates that partially corrupted structures are necessary to achieve better results. For some molecules (i.e., Ethanol and Salicycli acid), DeNS with noise added to all atoms (Index 5) can be worse than training without DeNS (Index 1).
{adjustwidth}
@@
-3.5cm-3.5cm Method Aspirin Benzene Ethanol Malonaldehyde Naphthalene Salicylic acid Toluene Uracil Force $\lambda_{E}\neq 0$ Partially corrupted Index DeNS encoding in Eq. 6 structures energy forces energy forces energy forces energy forces energy forces energy forces energy forces energy forces 1 5.3 7.2 2.2 6.6 2.2 3.1 3.3 5.8 3.7 2.1 4.5 4.1 3.8 2.1 4.3 3.3 2 5.1 5.7 2.3 6.1 2.2 2.6 3.2 4.4 3.7 1.7 4.4 3.7 3.5 1.9 4.2 3.3 3 8.6 9.1 2.3 6.3 2.3 3.3 3.2 5.8 7.7 6.1 5.2 10.6 3.7 2.0 5.5 6.5 4 5.1 5.8 2.3 6.1 2.3 2.6 3.2 4.5 3.7 1.7 4.8 3.6 3.7 1.9 4.2 3.3 5 5.3 6.9 2.4 6.3 2.3 3.2 3.3 5.4 3.7 2.0 4.8 4.2 4.0 2.0 4.2 3.8
@@
Table 6: Ablation study on the design choices of DeNS using MD17 dataset. Mean absolute error results are evaluated on the testing set. Energy and force are in units of meV and meV/Å. Index 1 and Index 2 correspond to “Equiformer ($L_{max}=2$)” and “Equiformer ($L_{max}=2$) + DeNS” in Table 4.
@@ 5 Conclusion
In this paper, we propose to use denoising non-equilibrium structures (DeNS) as an auxiliary task to better leverage training data and improve performance on original tasks of energy and force predictions. Denoising non-equilibrium structures can be an ill-posed problem since there are many possible target structures. To address the issue, we propose force encoding and take the forces of original structures as inputs to specify which non-equilibrium structures we are denoising. With force encoding, DeNS successfully improves the performance on original tasks when it is used as an auxiliary task. We conduct extensive experiments on OC20, OC22 and MD17 datasets to demonstrate that DeNS can boost the performance on energy and force predictions across datasets of various scales with minimal increase in training cost and is applicable to many equivariant networks.
@@ Appendix A Related Works
A.1 Comparison to Previous Works on Denoising
We discuss previous works on denoising (Godwin et al., 2022; Zaidi et al., 2023; Feng et al., 2023b; Wang et al., 2023) in chronological order and compare them with this work as below.
Godwin et al. (2022) first proposes the idea of adding noise to 3D coordinates and then using denoising as an auxiliary task. The auxiliary task is trained along with the original task without relying on another large dataset. Their approach requires known equilibrium structures and therefore is limited to QM9 (Ramakrishnan et al., 2014; Ruddigkeit et al., 2012) and OC20 IS2RE datasets and can not be applied to force prediction such as OC20 S2EF dataset. For QM9, all the structures are at equilibrium, and for OC20 IS2RE, the target of denoising is the relaxed, equilibrium structure. Denoising without force encoding is well-defined on both QM9 and OC20 IS2RE datasets. In contrast, this work proposes using force encoding to generalize their approach to non-equilibrium structures, which have much larger datasets than equilibrium ones. Force encoding can achieve better results on OC20 S2EF-2M dataset with little overhead (Index 2 and Index 3 in Table 1(b)) and is indispensable on MD17 dataset (Section 4.3).
Feng et al. (2023b) follows the same practice of pre-training via denoising (Zaidi et al., 2023) and proposes a different manner of adding noise. Specifically, they separate noise into dihedral angle noise and coordinate noise and only learn to predict coordinate noise. However, adding noise to dihedral angles requires tools like RDKit (rdk, ) to obtain rotatable bonds and cannot be applied to other datasets like OC20 and OC22.
Although Zaidi et al. (2023) and Feng et al. (2023b) report results of force prediction on MD17 dataset, they first pre-train models on PCQM4Mv2 dataset (Nakata & Shimazaki, 2017) and then fine-tune the pre-trained models on MD17 dataset. We note that their setting is different from ours since we do not use any dataset for pre-training. As for fine-tuning on MD17 dataset, Zaidi et al. (2023) simply follows the same practice in standard supervised training. Feng et al. (2023b) explores fine-tuning with objectives similar to Noisy Nodes (Godwin et al., 2022), but the performance gain is much smaller than ours. Concretely, in Table 5 in Feng et al. (2023b), the improvement in force prediction on Aspirin is about $2.6\%$ while we improve force MAE by $20.8\%$.
A.2 SE(3)/E(3)-Equivariant Networks
@@ B.1 Training Details
Since each structure in OC20 S2EF dataset has a pre-defined set of fixed and free atoms and we only predict forces of free atoms, we only apply DeNS to free atoms. When partially corrupted structures are used, we add noise to and denoise a random subset of free atoms. When training EquiformerV2 on OC20 S2EF-All+MD dataset, we only apply DeNS to structures from the All split.
We add an additional block of equivariant graph attention to EquiformerV2 for noise prediction. We mainly follow the hyper-parameters of training EquiformerV2 without DeNS on OC20 S2EF-2M and S2EF-All+MD datasets. For training EquiformerV2 on OC20 S2EF-All+MD dataset, we increase the number of epochs from $1$ to $2$ for better performance. This results in higher training time than other methods. However, we note that we already demonstrate training with DeNS can achieve better results given the same amount of training time in Table 1(a). Table 7 summarizes the hyper-parameters of training EquiformerV2 with DeNS for the ablation studies on OC20 S2EF-2M dataset in Section 4.1.1 and for the main results on OC20 S2EF-All+MD dataset in Section 4.1.2. Please refer to the work of EquiformerV2 (Liao et al., 2023) for details of the architecture. For training eSCN with DeNS, we use the same DeNS-related hyper-parameters as those for training EquiformerV2 for $20$ epochs and the same module as force prediction to predict noise.
V100 GPUs with 32GB are used to train models. We use $16$ GPUs for training EquiformerV2 for 12 epochs and eSCN on OC20 S2EF-2M dataset and use $32$ GPUs for training EquiformerV2 for 20 and 30 epochs. We train EquiformerV2 with $128$ GPUs on OC20 S2EF-All+MD dataset. The training time and the numbers of parameters of different models on OC20 S2EF-2M dataset can be found in Table 1(a). For OC20 S2EF-All+MD dataset, the training time is $88982$ GPU-hours and the number of parameters is $160$M.
@@ C.1 Training Details
Different from OC20, all the atoms in a structure in OC22 are free, and we apply DeNS to all the free atoms. We add an additional block of equivariant graph attention to EquiformerV2 for noise prediction. We follow the same hyper-parameters not relevant to DeNS, and Table 8 summarizes the hyper-parameters for the results on OC22 in Table 3. We use $32$ V100 GPUs (32GB) for training. The training time is $5082$ GPU-hours, and the number of parameters $127$M.
@@
Hyper-parameters OC20 S2EF-2M dataset OC20 S2EF-All+MD dataset
@@
Batch size $64$ for $12$ epochs $512$
$128$ for $20,30$ epochs
@@
Probability of optimizing DeNS $p_{\text{DeNS\leavevmode\nobreak\ }}$ $0.5$ $0.125$
DeNS coefficient $\lambda_{\text{DeNS\leavevmode\nobreak\ }}$ $10$ $15$
Standard deviation of Gaussian noise $\sigma$ $0.1$ for $12$ epochs $0.1$
$0.15$ for $20,30$ epochs
Corruption ratio $r_{\text{DeNS}}$ $0.5$ $0.25$
Table 7: Hyper-parameters of training EquiformerV2 with DeNS on OC20 S2EF-2M dataset and OC20 S2EF-All+MD dataset.
@@
DeNS coefficient $\lambda_{\text{DeNS\leavevmode\nobreak\ }}$ $25$
Standard deviation of Gaussian noise $\sigma$ $0.15$
@@ D.2 Training Details
We use the official implementation of Equiformer (Liao & Smidt, 2023) for experiments on the MD17 dataset and follow most of the original hyper-parameters for training with DeNS. For training DeNS, we use an additional block of equivariant graph attention for noise prediction, which slightly increases training time and the number of parameters. The hyper-parameters introduced by training DeNS and the values of energy coefficient $\lambda_{E}$ and force coefficient $\lambda_{F}$ on different molecules can be found in Table 9. Empirically, we find that linearly decaying DeNS coefficient $\lambda_{\text{DeNS}}$ to 00 thoughout the training can result in better performance. For the Equiformer variant without attention and layer normalization, we find that using normal distributions to initialize weights can result in training divergence and therefore we use uniform distributions. For some molecules, we find training Equiformer variant without attention and layer normalization with DeNS is unstable and therefore reduce the learning rate to $3\times 10^{-4}$.
@@
{adjustwidth}
-2.5cm-2.5cm Hyper-parameter Aspirin Benzene Ethanol Malonaldehyde Naphthalene Salicylic acid Toluene Uracil Energy coefficient $\lambda_{E}$ 1 1 1 1 2 1 1 1 Force coefficient $\lambda_{F}$ 80 80 80 100 20 80 80 20 Probability of optimizing DeNS $p_{\text{DeNS}}$ 0.25 0.25 0.25 0.25 0.25 0.25 0.125 0.25 DeNS coefficient $\lambda_{\text{DeNS}}$ 5 5 5 5 5 5 5 5 Standard deviation of Gaussian noises $\sigma$ 0.05 0.05 0.05 0.05 0.05 0.05 0.025 0.05 Corruption ratio $r_{\text{DeNS}}$ 0.25 0.25 0.25 0.25 0.25 0.25 0.25 0.25
Table 9: Hyper-parameters of training Equiformer ($L_{max}=2$) and Equiformer ($L_{max}=3$) with DeNS on the MD17 dataset. Other hyper-parameters not listed here are the same as the original Equiformer trained without DeNS.